How many infrared and Raman absorptions do you expect for BF3 if it had each of the following geometries. show work

(a) Planar (b) Pyramidal (c) T-shaped
Answer #1

BF3 molecule has total 12 degrees of freedom In general, we would expect 3N-6 = 3x4-6=6 degrees of freedom. However, is the m- Now, to has 6 for Non-symmetric, like fundamental frequency T shape as general. . No. of Infrared and Raman absorption in pSymmetric planar Active modes 04-40 Ramen = active Infrared = inactive 80 Raman - inactive Infrared = active ④ = upwards 0 =2. No of Infrared and famen active in Pyramidal 1. Bf3 (Point group (31) All distinct 4 fundamental vibration frequenze are t1. 3. No of infrared and Raman active more in T-shaped Bf3 ( point group. Ev.) All distinct 6 fundamental frequencies are botActive modes Non-symmetric T shaped f-B-F All modes are FL both Raman and Infrared active. ►FB- Er F-B-f6 = upwords - down words -